OpenStudy (anonymous):

A freshly brewed cup of coffee has temperature 95°C in a 20°C room. When its temperature is 79°C, it is cooling at a rate of 1°C per minute. When does this occur? (Round your answer to two decimal places.)

OpenStudy (anonymous):

This is a rate of change problem, so you are going to need calculus - we need some assumptions here - as to the type of relationship. At the moment, you know dT/dt at a particular time, and To

OpenStudy (anonymous):

Wouldn't it be something along the lines of dT/dt = 20+95e^kt?

OpenStudy (anonymous):

Then we would go ahead and substitute for k, giving us T(79) = 20+95e^(1)k?

OpenStudy (anonymous):

I mean 20+95e^1t, substituting for k

OpenStudy (anonymous):

ok, so you are modelling temp using the exponential function, and k will be negative. Let T=20+75e^(kt) so when t=0, T=95 (since 20 +75=95). When T=79 we have 79=20+75e^(kt) (1) and differentiating to get dT/dt=k*75e^(kt)=-1 (2) So We have two equations, rearrange (1) to get what e^kt equals substitute it into the second to solve for k, then you can solve for t using either (1) or (2) and ln

OpenStudy (anonymous):

You can solve this now?